OpenStudy (anonymous):

Please Weights of male mountain lions follow the normal distribution with a median of 150 lb and an interquartile range of 8.2 lb. Find the 75th percentile of the weights. [Calculate your answer in pounds (lb), but do not enter the units of measurement.]

OpenStudy (kropot72):

The standard deviation can be calculated using the formula \[\sigma=\frac{Interquartile\ range}{1.34896}\] In this case the standard deviation 8.2/1.34896= 6.079 A standard normal distribution table shows the z-score for a cumulative probability of 75% is z = 0.675 The median and the mean are the same. Therefore the following formula can be used: \[z=\frac{X-\mu}{\sigma}\] Substituting we get \[0.675=\frac{X-150}{6.079}\] Solving for X gives X = 150 + 4.1